A company that is producing cork pens has introduced a new machine that reduces processing time for each unit; output is increased by 25 percent with less material, but one additional worker is required. For the new machine, material is now $10/unit. Overhead is charged at 1.6 times direct labor cost. Finished units sell for $31 each. Under the old process, five workers could produce in total 60 units per hour. Labor costs are $12/hour, and material input was previously $16/unit. For the new machine, material is now $10/unit. Overhead is charged at 1.6 times direct labor cost. Finished units sell for $31 each.
1.1. Calculate the multifactor productivity before and after the process improvement.
1.2. Calculate how much was the increase in productivity is associated with the process improvement?
1.3. List two ways how this company can increase more the productivity without increasing the costs.
Answer #1

1.1

MFP before process improvement = 60/(labor cost + overhead cost + material cost)

MFP before process improvement = 60/(5*12 + 1.6*5*12 + 16*60)

MFP before process improvement = .05376 units per dollar spent

MFP after process improvement = 60*1.25/(labor cost + overhead cost + material cost)

MFP after process improvement = 60*1.25/(6*12 + 1.6*6*12 + 10*60*1.25)

MFP after process improvement = .08003 units per dollar spent

==

1.2

Increase in productivity due to process improvement = (.08003 - . 05376)/. 05376

Increase in productivity due to process improvement = 48.87%

==

1.3

The first way is to given department training using internal trainers so that they perform their task with higher efficiency. It will increase productivity. The second way is to convince workers, operate in second shift so that more number of hours of work is contributed and productivity increases.